# MABViT - Modified Attention Block Enhances Vision Transformers

### 03 Dec 2023 v1

@@ Introduction
The challenge with the representation capability of Pre-LN Transformers was first identified by Admin (Liu et al. 2020). As the number of layers increases, the $X$ term grows linearly, making the output from the Multi-Head Attention, or MLP blocks, relatively insignificant.
@@
In deeper layers:
@@
If $X_{i}\approx X_{j}$, this implies that $X_{i+2}\approx X_{j+1}$, supporting our hypothesis that the linear growth of the $X$ term causes similarity in performance between parallel and standard Transformer formulations.
In Post-LN Transformers, the normalization of the $X$ term takes place after every Multihead-attention and MLP block, a practice associated with training instability. Admin noted that when Post-LN transformers do not diverge, they tend to outperform the Pre-LN variant in machine translation. However, in the context of most vision tasks, Post-LN transformers often exhibit divergence, motivating us to develop a new transformer variant aimed at surpassing the performance of traditional Vision Transformers.
@@
In Pre-LN Transformers, theory suggests that incorporating the Multi-Head Attention output into the residual before feeding it to the MLP block becomes redundant in deeper layers. To address this, we propose integrating the activation function within the attention module to apply a non-linear transformation to the Multi-Head Attention output effectively.
@@
There is a noticeable difference between the validation accuracy of the standard and parallel architectures across all the models. With smaller-scale models, the magnitude of the $X$ term is comparable to the output of the Multi-Head Attention block. Consequently, the standard architectures attain superior performance compared to the parallel formulations. We observe that the difference in results reduces when we increase the width of the layers. We hypothesize that the magnitude of the $X$ term rises with greater width, causing the parallel and standard structures to perform more similarly.

### 01 Jan 2024 v2

@@ Introduction
The challenge with the representation capability of Pre-LN transformers was first identified by Admin (Liu et al. 2020). As the number of layers increases, the $X$ term grows, making the output from the Multi-Head Attention, or MLP blocks, relatively insignificant.
@@
Considering an initial 1x1 tensor as input and positive outputs from both the Multihead-Attention and MLP blocks at each layer, as we progress through deeper layers:
@@
If $X_{i}\approx X_{j}$, it implies that $X_{i+2}\approx X_{j+1}$, supporting our hypothesis that the growth of the $X$ term causes similarity in performance between parallel and standard transformer formulations under certain cases.
In Post-LN transformers, the normalization of the $X$ term takes place after every Multihead-attention and MLP block, a practice associated with training instability. Admin (Liu et al. 2020) noted that when Post-LN transformers do not diverge, they tend to outperform the Pre-LN variant in machine translation. However, in the context of most vision tasks, Post-LN transformers often exhibit divergence, motivating us to develop a new transformer variant aimed at surpassing the performance of traditional ViTs.
@@
In Pre-LN transformers, we hypothesize that incorporating the Multi-Head Attention output into the residual before feeding it to the MLP block becomes redundant in deeper layers. To address this, we propose integrating the activation function within the attention module to apply a non-linear transformation to the Multi-Head Attention output effectively.
@@
There is a noticeable difference between the validation accuracy of the standard and parallel architectures across all the models. With smaller-scale models, the magnitude and variance of the $X$ term are reduced. Consequently, the standard architectures attain superior performance compared to the parallel formulations. We observe that the difference in results reduces when we increase the width of the layers. We hypothesize that the magnitude of the $X$ term rises with greater width, causing the parallel and standard structures to perform more similarly.
@@ Conclusion
Acknowledgement
We extend our sincere thanks to Bharat Kumar Sharma and NVIDIA for their crucial support in the project.